OpenStudy (anonymous):

True or false. Linear Algebra. If A is invertible and 1 is an eigenvalue for A, then 1 is also an eigenvalue of A^-1.

jimthompson5910 (jim_thompson5910):

If 1 is an eigenvalue for the matrix A, then Ax = 1*x Ax = x A^(-1)*Ax = A^(-1)x x = A^(-1)x A^(-1)x = x A^(-1)x = 1*x So 1 is also an eigenvalue of A^(-1)

OpenStudy (amistre64):

nice use of definition